OpenStudy (anonymous):

2x - 2y + 5z - 2x - y + 3z

OpenStudy (anonymous):

do you know what like terms are ?

OpenStudy (anonymous):

Simplify the following expression:

OpenStudy (mathmale):

Please always include the instructions with each posting such as this one. Building on kellie's question: To "simplify the following expression," combine "liike terms." Provided that you know what "like terms" are, that's all you have to do.

OpenStudy (anonymous):

no the answer

OpenStudy (mathmale):

What do you mean by "no the answer"? No what or know what? 1) explain in your own words what "like terms" mean. 2) combine the like terms in the given expression.

OpenStudy (anonymous):

add your like terms.. 2x - 2x = -2y - y = 5z + 3z = Then put it all together...you do the math and what do you get ?

OpenStudy (anonymous):

you gonna at least try ?

OpenStudy (anonymous):

-3y+what?

OpenStudy (mathmale):

No more comments until @williamsc258 responds, OK? Thank you.